# Proxy-Tuning: Tailoring Multimodal Autoregressive Models for Subject-Driven Image Generation

### v1

@@ 6 Conclusion
This paper addresses subject-driven image generation using autoregressive (AR) models. While direct fine-tuning AR models fails to capture subject-specific characteristics, our Proxy-Tuning leverages diffusion models as supervisors to enable effective personalization. Beyond demonstrating robustness across various training configurations, our method reveals a surprising weak-to-strong generalization phenomenon: AR models consistently surpass their diffusion supervisors in both subject fidelity and prompt adherence. Furthermore, AR models exhibit superior parameter efficiency by learning multiple subjects simultaneously, where diffusion models struggle. These advantages establish AR models, when properly guided through Proxy-Tuning, as a scalable and powerful solution for widespread multimodal applications.
7 Limitation
While Proxy-Tuning enables strong subject-driven image generation capability for AR models, several aspects remain to be explored. First, a theoretical understanding of the weak-to-strong phenomenon deserves deeper investigation, which could provide insights into the fundamental advantages of AR architectures in image generation. Second, our current approach requires empirical selection of proxy dataset sizes; developing automated strategies for determining optimal proxy data quantity could further improve the method’s efficiency and accessibility.
@@ 7 Limitation
Appendix A User Study Details
We conduct user study to better validate the qualitative performance of our Proxy-Tuning. The study involves 9 participants. The participants are asked to grade the subject images based on three criteria: image quality, subject fidelity, and prompt fidelity. A comprehensive grading is required for the several groups of images of the four methods in the context of three cases. Each case is composed of a subject image and several groups of images obtained from four methods. In each group, there is one piece of text and two images. The results are rated on a scale of 1 to 5, with 5 being the highest. We include direct subject tuning of diffusion and AR models, Proxy-Tuning of diffusion models, and our proposed Proxy-Tuning of AR models.
Appendix B Training Details
The training details are illustrated in Table 8. We follow the license and intended use of these open-source methods.
Learning Rate Batch Size Optimizer Training Steps Model Size
SDXL 1e-4 4 AdamW 500 2.6B
SD3 Medium 4e-4 4 AdamW 500 2B
SD3.5 Large 4e-4 4 AdamW 500 8B
FLUX.1 [dev] 1.0 4 Prodigy 500 12B
LlamaGen-XL 1e-4 8 AdamW 1250 775M
Lumina-mGPT 2e-5 4 AdamW 2500 7B
Table 8: Training details of different models.

### v2

@@ 6 Conclusion and Limitations
This paper addresses subject-driven image generation using autoregressive (AR) models. While direct fine-tuning AR models fails to capture subject-specific characteristics, our Proxy-Tuning leverages diffusion models as supervisors to enable effective personalization. Beyond demonstrating robustness across various training configurations, our method reveals a surprising weak-to-strong generalization phenomenon: AR models consistently surpass their diffusion supervisors in both subject fidelity and prompt adherence. Furthermore, AR models exhibit superior parameter efficiency by learning multiple subjects simultaneously, where diffusion models struggle. These advantages establish AR models, when properly guided through Proxy-Tuning, as a scalable and powerful solution for widespread multimodal applications. However, several aspects remain to be explored. First, a theoretical understanding of the weak-to-strong phenomenon deserves deeper investigation, which could provide insights into the fundamental advantages of AR architectures in image generation. Second, our current approach requires empirical selection of proxy dataset sizes; developing automated strategies for determining optimal proxy data quantity could further improve the method’s efficiency.